OpenStudy (anonymous):

Not sure if I am doing this right..... 14/6y+11/9y

OpenStudy (amistre64):

is y under the fractions?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

then i assume it wants you to add these together. Assume that there is some value which we will call D out there that this is equal to. \[\frac{14}{6y}+\frac{11}{9y} = D\] multiply it all by the lowest denom \[\frac{14(6y)}{6y}+\frac{11(6y)}{9y} = (6y)D\] and simplify \[14+\frac{11(2)}{3} = 6yD\] now multiply it all again by the lowest denom and simplify \[14(3)+\frac{11(2)(3)}{3} = 6y(3)D\] \[14(3)+11(2) = 18yD\] \[42+22 = 18yD\] \[42+22 = 18yD\] i notice here that they are all even so i can divide out a 2 \[42/2+22/2 = 18yD/2\] \[21+11 = 9yD\] \[32 = 9yD\] and finally divide off that 9y \[\frac{32}{9y} = \frac{9yD}{9y}\] \[\frac{32}{9y} = D\] and that should be the brunt of it

OpenStudy (amistre64):

of course there is simpler ways to go about it, but they have to deal with remembering certain tricks instead of letting the math work it out

OpenStudy (anonymous):

Thank u for the explanation!! I needed that. :)

OpenStudy (amistre64):

yer welcome ;)